OpenStudy (alivejeremy):

Efrain must spend less than $76 on his phone bill this month. He pays a monthly fee of $19, plus an additional $3 for each long-distance call. 3x + 19 < 76 How may long-distance calls can Efrain make this month? x < 57 x < 38 x < 3 x < 19

OpenStudy (alivejeremy):

help please

OpenStudy (alivejeremy):

!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (alivejeremy):

can anybody help

OpenStudy (alivejeremy):

can anybody help me

OpenStudy (alivejeremy):

can anybody help me

OpenStudy (alivejeremy):

u bk den

OpenStudy (yummy123):

yes

OpenStudy (alivejeremy):

Efrain must spend less than $76 on his phone bill this month. He pays a monthly fee of $19, plus an additional $3 for each long-distance call. 3x + 19 < 76 How may long-distance calls can Efrain make this month? x < 57 x < 38 x < 3 x < 19

OpenStudy (alivejeremy):

can someone help

OpenStudy (yummy123):

is this on studyisland

OpenStudy (alivejeremy):

yea

OpenStudy (welshfella):

solve for x first subtract 19 from both sides of the < then divide both sides by 3 to give x < something

OpenStudy (alivejeremy):

?

OpenStudy (alivejeremy):

?

OpenStudy (welshfella):

3x + 19- 19 < 76 - 19 can you simplify that?

OpenStudy (welshfella):

what is 19-19?

OpenStudy (anonymous):

You can multiply each of the numbers that you have for each choice with 3. Then add 19 and see if it is less than 76.

OpenStudy (alivejeremy):

19-19=0

OpenStudy (welshfella):

right so 3x = 76 - 19 what is 76-19?

OpenStudy (alivejeremy):

57

OpenStudy (alivejeremy):

Gina is having a party and wants to spend less than $480. She has already spent $400. The only item left on her list is pizzas, which cost $10 each. How many pizzas, x, can she purchase and stay under her budget? x < 40 x < 8 x < 48 x < 70

OpenStudy (welshfella):

so the inequality becomes 3x < 57 to get x < you now divide both sides by 3

OpenStudy (welshfella):

3x / 3 = x and 57/3 = ?

OpenStudy (welshfella):

??

OpenStudy (alivejeremy):

bk

OpenStudy (alivejeremy):

u still there

OpenStudy (welshfella):

yea

OpenStudy (alivejeremy):

ok

OpenStudy (alivejeremy):

Gina is having a party and wants to spend less than $480. She has already spent $400. The only item left on her list is pizzas, which cost $10 each. How many pizzas, x, can she purchase and stay under her budget? x < 40 x < 8 x < 48 x < 70

OpenStudy (welshfella):

what about the first question what is the answer to that one?

OpenStudy (alivejeremy):

tht is the first question

OpenStudy (welshfella):

I@m talking about the question regarding long-distance calls.

OpenStudy (alivejeremy):

oh i got tht right

OpenStudy (alivejeremy):

lol

OpenStudy (alivejeremy):

bk

OpenStudy (alivejeremy):

Carlene is selling T-shirts for a fundraiser. The T-shirts cost $20 each. To date, she has raised $700. Her goal is to raise more than $8,700. How many more T-shirts, x, does Carlene need to sell to reach her goal?

OpenStudy (alivejeremy):

nvm tht one

OpenStudy (mathmale):

Please post just one question at a time, OK? Otherwise other users may not know what to focus on, and time may be lost. Please repost your "Carlene" question.

OpenStudy (welshfella):

10x + 400 < 480 - solve this to get answer to the pizza question

OpenStudy (alivejeremy):

ok ima make a new question